OpenStudy (anonymous):

slope that pass (-6,1) (4,-4)

OpenStudy (anonymous):

To find the slope you would use this formula....(Y2-Y1)/(X2-X1)

OpenStudy (hba):

Use the formula , \[\huge\ m=\frac{ y_2-y_1 }{ x_2-x_1 }\] Here (x1,y1)=(-6,1) (x2,y2)= (4,-4)

OpenStudy (anonymous):